OpenStudy (jordanloveangel):

In May, Liam and Charlie had the same amount of money in their savings accounts. In June, Liam deposited $140 into his account. Charlie said he increased the money in his account by 5%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in May? $?

OpenStudy (jordanloveangel):

@linn99123

OpenStudy (jordanloveangel):

@sammie2017

OpenStudy (jordanloveangel):

@RyleeC

OpenStudy (anonymous):

m+140=x m+.05m=x

OpenStudy (anonymous):

so that m+.05m=m+140

OpenStudy (anonymous):

m=2800

OpenStudy (jordanloveangel):

so the answer is $2940

OpenStudy (anonymous):

m = money in may

OpenStudy (anonymous):

answer is 2800

OpenStudy (anonymous):

in june they had 2940

OpenStudy (jordanloveangel):

1. X + 140 = X + (X x 0.05) in June 2. 140 = X x 0.05 3. 140/0.05 = X 4. X = 2800

OpenStudy (jordanloveangel):

like this right?

OpenStudy (anonymous):

I think use * as multiply